# Entanglement phases and phase transitions in monitored free fermion system due to localizations

### v1

@@ Abstract
In recent years, the presence of local potentials has significantly enriched and diversified the entanglement patterns in monitored free fermion systems. In our approach, we employ the stochastic Schrödinger equation to simulate a one-dimensional spinless fermion system under continuous measurement and local potentials. By averaging the steady-state entanglement entropy over many quantum trajectories, we investigate its dependence on measurement and localization parameters. We used a phenomenological model to interpret the numerical results, and the results show that the introduction of local potentials does not destroy the universality class of the entanglement phase transition, and that the phase boundary is jointly characterized by the measurement process and the localization mechanism. This work offers a new perspective on the characterization of the entanglement phase boundary arising from the combined effects of measurement and localization, and provides criteria for detecting this novel phase transition in cold atom systems, trapped ions, and quantum dot arrays.
@@ I Introduction
Classical paradigms of nonequilibrium quantum physics-such as thermalization under the eigenstate thermalization hypothesis [1, 2, 3, 4, 5] and its counterpoint in the quantum Zeno effect [6, 7, 8, 9, 10], offer a striking tension: while coherent dynamics generically scramble information and build volume-law entanglement, frequent observations can freeze evolution by continually projecting local degrees of freedom. This tension naturally leads to the notion of a measurement-driven reorganization of quantum states: as the rate or strength of monitoring is increased, the balance between entanglement production and removal can qualitatively change, suggesting a sharp transition in the long-time entanglement structure. This phenomenon is now understood as a measurement-induced phase transition (MIPT), in which entanglement scaling morphs from volume law to area law due to the backaction of measurements.
A broad body of work has established both the ubiquity and diversity of MIPTs across platforms and protocols [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. In random quantum circuits [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], interleaving unitary gates with stochastic measurements yields a finite-threshold transition. Related transitions appear under Hamiltonian dynamics with projective or continuous monitoring [44, 45, 46, 47, 48], and persist for both projective and weak measurements. Across these settings, the steady-state phases (volume, logarithmic/sub-volume, and area) and their critical properties depend on symmetries [41, 52, 53, 12], dimensionality  [54], and the microscopic structure of measurements [30, 55, 56, 57, 58, 59], revealing a rich taxonomy of measurement-driven universality classes.
@@
In the presence of localization, the landscape becomes more intricate. The root cause lies in the fact that localization tends to confine particles or quantum information to local regions, thereby significantly suppressing the generation and propagation of long-range entanglement. Different types of localization affect entanglement and information spreading through their respective physical mechanisms, and may thus alter the stability of phases and critical behavior in different ways. For example, disorder or quasi-periodic potential(QPP) induced Anderson localization and Stark potential(SP) induced Stark localization imprint distinct single-particle spectra, spatial profiles and transport properties that can [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81], in principle, reshape measurement-driven entanglement phases and transitions.
@@
$H=-J\sum_{j}(c_{j}^{\dagger}c_{j+1}+\mathrm{h.c.})+\sum_{j}P_{j}\,n_{j},$ (1)
where $c_{j}$ and $c^{\dagger}_{j}$ are the annihilation and creation operators of the spinless fermion at site $j$, $P_{j}$ represents the strength of potential at lattice site $j$. We initialize the system in a Néel state, that is, $\psi(t=0)=$ $\prod_{i=1}^{L/2}c_{2i-1}^{\dagger}|\mathrm{vac}\rangle$, where $|\mathrm{vac}\rangle$ is the vacuum state, and consider its evolution under open boundary conditions (OBCs). We perform a quantum measurement uniformly for all sites with a finite measurement amplitude (free quency) $\gamma$:
@@ III.1 Phase boundaries
To validate Eq. (11), we first map the effective central charge $\bar{c}_{\mathrm{eff}}$ over the parameter space spanned by the measurement strength $\gamma$ and the local potential strength (either Stark, $\Delta$, or quasi-periodic, $V$), thereby characterizing the scaling regimes of the entanglement entropy. As shown in Fig. 2, the phase diagrams under varying measurement and potential strengths are presented for the SP and QPP systems. The colormap encodes $\bar{c}_{\mathrm{eff}}$ extracted from half-chain entanglement entropy; the horizontal axis is the measurement rate $\gamma$, and the vertical axis is the potential strength $\Delta$ or $V$. The diagrams clearly separate area-law and log-law regimes: regions with large $\bar{c}_{\mathrm{eff}}$ correspond to log-law growth (an entanglement-spreading phase), whereas regions with $\bar{c}_{\mathrm{eff}}\approx 0$ exhibit suppressed entanglement.
@@
We find that increasing either the SP or QPP shrinks the entangled region and eventually eliminates the phase boundary, although the deformation is qualitatively different between SP and QPP. The dashed curves are fits to Eq. (11), which agree closely with critical points obtained via finite-size scaling (black dots). As the potential strength increases, the entangled region narrows, the boundary bends, and eventually terminates, reflecting the progressively stronger suppression of entanglement propagation caused by localization. Importantly, the boundary deforms differently in the two cases: for SP, it changes super-exponentially with $\Delta^{2}$; for QPP, it changes exponentially with $V$ results that are fully consistent with the analytical structure established earlier. Together, they reveal a measurement-localization coupled entanglement phase transition: when the localization potential is small, the phase boundary is primarily determined by the measurement strength, and its descent is relatively gradual; as the localization potential increases, its suppression of entanglement growth causes the boundary to drop rapidly. However, because the Stark localization induced by SP is stronger than the Anderson localization induced by QPP, the phase boundary in Fig. 2(a) falls faster, reaching the area-law region much earlier.
@@ III.2 Phases and phase transitions
We next follow the finite-size flow of the effective central charge $\bar{c}_{\mathrm{eff}}$ under OBCs. For each size $L$ we fit the entanglement profile to Eq. (12) and obtain a single $\bar{c}_{\mathrm{eff}}$ subjected to different measurement intensities and local potentials, as illustrated in Fig. 4. Near criticality, the half-chain entanglement entropy exhibits logarithmic scaling consistent with non-unitary conformal field theory:
@@
where $\ell$ is the subsystem size, $\bar{c}_{\mathrm{eff}}$ is the effective central charge, and $s_{0}$ is a non-universal constant. In the thermodynamic limit, $\bar{c}_{\mathrm{eff}}\to 0$ in the area-law phase, while it remains finite in the logarithmic phase, thus serving as a diagnostic for the transition. As shown in Fig. 4(a,b), when the local potential is fixed at a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ causes $\bar{c}_{\mathrm{eff}}$ curves for different system sizes to gradually intersect, indicating a transition near $\gamma\approx 0.3$. The insets strongly corroborate this behavior: as $\gamma$ increases, the entanglement entropy transitions from rapid growth to saturation, signaling a crossover from log-law to area-law scaling. Similarly, in Fig. 4(c,d), when the measurement strength is fixed at a low value ($\gamma=0.1$), increasing the local potential strength leads to similar crossing behavior among different system sizes, with transition points occurring around $\Delta\approx 0.3$ and $V\approx 1.8$, respectively. The insets in Fig. 4 clearly demonstrate this effect: the entanglement entropy drops from steep growth to a nearly constant value as localization increases, again signifying a transition from log-law to area-law behavior. These results collectively indicate that both measurement strength and local potential can induce a phase transition in the system. In both cases, entanglement growth is suppressed, although the mechanisms and suppression profiles differ between SP and QPP. These observations reveal how both measurement and localization reduce boundary coupling and suppress entanglement growth. The distinct functional behaviors of SP and QPP reinforce the validity of $\bar{c}_{\mathrm{eff}}$ as a transition indicator.
@@ Appendix A Discussion on the Phase Boundary
this is the entanglement entropy growth rate corrected by the SP. Similarly, for a one-dimensional QPP,
$P_{j}\;=\;V\cos\bigl{(}2\pi\beta j+\varphi\bigr{)}\,$, (33)
$V$ is the strength of the QPP, the single-particle eigenfunctions localize exponentially with localization length $\xi(V)$. Hence the entanglement entropy growth rate is suppressed to

### v2

# Entanglement phases and phase transitions in monitored free fermion systems of localization

@@ Abstract
In recent years, the presence of local potentials has significantly enriched and diversified the entanglement patterns in monitored free fermion systems. In our approach, we employ the stochastic Schrödinger equation to simulate a one-dimensional spinless fermion system under continuous measurement and local potentials. By averaging the steady-state entanglement entropy over many quantum trajectories, we investigate its dependence on measurement and localization parameters. We used a phenomenological model to interpret the numerical results, show that the introduction of local potentials does not destroy the universality class of the entanglement phase transition, and the phase boundary is jointly characterized by the measurement process and the localization mechanism. This work offers a new perspective on the characterization of the entanglement phase boundary arising from the combined effects of measurement and localization, and provides criteria for detecting this novel phase transition in cold atom systems, trapped ions, and quantum dot arrays.
@@ I Introduction
Classical paradigms of nonequilibrium quantum physics-such as thermalization under the eigenstate thermalization hypothesis [1, 2, 3, 4, 5] and its counterpoint in the quantum Zeno effect [6, 7, 8, 9, 10], offer a striking tension: while coherent dynamics generically scramble information and build volume-law entanglement, frequent observations can freeze evolution by continually projecting local degrees of freedom. This tension naturally leads to the notion of a measurement-driven reorganization of quantum states: as the rate or strength of monitoring is increased, the balance between entanglement production and removal can qualitatively change, suggesting a sharp transition in the long-time entanglement structure. This phenomenon is now understood as a measurement-induced phase transition (MIPT) [11, 12, 13, 14, 15, 16, 17], in which entanglement scaling morphs from volume law to area law due to the backaction of measurements.
A broad body of work has established both the ubiquity and diversity of MIPTs across platforms and protocols [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In random quantum circuits [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], interleaving unitary gates with stochastic measurements yields a finite-threshold transition. Related transitions appear under Hamiltonian dynamics with projective or continuous monitoring [51, 52, 53, 54, 55], and persist for both projective and weak measurements. Across these settings, the steady-state phases (volume, logarithmic/sub-volume, and area) and their critical properties depend on symmetries [48, 58, 59, 19], dimensionality  [60], and the microscopic structure of measurements [37, 61, 62, 63, 64, 65], revealing a rich taxonomy of measurement-driven universality classes.
@@
In the presence of localization, the landscape becomes more intricate. The root cause lies in the fact that localization tends to confine particles or quantum information to local regions, thereby significantly suppressing the generation and propagation of long-range entanglement. Different types of localization affect entanglement and information spreading through their respective physical mechanisms, and may thus alter the stability of phases and critical behavior in different ways. For example, disorder or quasi-periodic potential (QPP) induced Anderson localization and Stark potential (SP) induced Stark localization [66, 67, 68] imprint distinct single-particle spectra, spatial profiles and transport properties that can [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], in principle, reshape measurement-driven entanglement phases and transitions.
@@
$H=-J\sum_{j}\left(c_{j}^{\dagger}c_{j+1}+\mathrm{h.c.}\right)+\sum_{j}P_{j}\,n_{j},$ (1)
$P_{j}$ represents the strength of potential at lattice site $j$:
$P_{j}=\begin{cases}\Delta j/L,&\text{SP},\\ V\cos\left(\dfrac{2\pi j}{\tau}+\theta\right),&\text{QPP},\end{cases}$ (2)
where $J$ is the strength of the nearest-neighbor hopping term, $c_{j}$ and $c^{\dagger}_{j}$ are the annihilation and creation operators of the spinless fermion at site $j$. Here, $\Delta$ denotes the SP strength, $L$ is the number of lattice sites in the system, and the parameters $V$, AA, and $B$ represent, respectively, the QPP strength, the golden mean $(\sqrt{5}+1)/2$, and the phase shift of the potential. We initialize the system in a Néel state, that is, $\psi(t=0)=$ $\prod_{i=1}^{L/2}c_{2i-1}^{\dagger}|\mathrm{vac}\rangle$, where $|\mathrm{vac}\rangle$ is the vacuum state, and consider its evolution under open boundary conditions (OBCs). We perform a quantum measurement uniformly for all sites with a finite measurement amplitude (free quency) $\gamma$:
@@ III.1 Phase boundaries
Eq.(12) provides a phenomenological envelope-function expression [73], and in the Anderson case [102], the measurement further introduces a super-exponential envelope that strongly suppresses the coefficient $C_{0}$. To validate Eq. (12), we first map the effective central charge $\bar{c}_{\mathrm{eff}}$ over the parameter space spanned by the measurement strength $\gamma$ and the local potential strength (either Stark, $\Delta$, or quasi-periodic, $V$), thereby characterizing the scaling regimes of the entanglement entropy. As shown in Fig. 2, the phase diagrams under varying measurement and potential strengths are presented for the SP and QPP systems. The colormap encodes $\bar{c}_{\mathrm{eff}}$ extracted from half-chain entanglement entropy; the horizontal axis is the measurement rate $\gamma$, and the vertical axis is the potential strength $\Delta$ or $V$. The diagrams clearly separate area-law and log-law regimes: regions with large $\bar{c}_{\mathrm{eff}}$ correspond to log-law growth (an entanglement-spreading phase), whereas regions with $\bar{c}_{\mathrm{eff}}\approx 0$ exhibit suppressed entanglement.
@@
We find that increasing either the SP or QPP shrinks the entangled region and eventually eliminates the phase boundary, although the deformation is qualitatively different between SP and QPP. The dashed curves are fits to Eq. (12), which agree closely with critical points obtained via finite-size scaling (black dots). As the potential strength increases, the entangled region narrows, the boundary bends, and eventually terminates, reflecting the progressively stronger suppression of entanglement propagation caused by localization. Importantly, the boundary deforms differently in the two cases: for SP, it changes super-exponentially with $\Delta^{2}$; for QPP, it changes exponentially with $V$ results that are fully consistent with the analytical structure established earlier. Together, they reveal a measurement-localization coupled entanglement phase transition: when the localization potential is small, the phase boundary is primarily determined by the measurement strength, and its descent is relatively gradual; as the localization potential increases, its suppression of entanglement growth causes the boundary to drop rapidly. However, because the Stark localization induced by SP is stronger than the Anderson localization induced by QPP, the phase boundary in Fig. 2(a) falls faster, reaching the area-law region much earlier. It is worth noting that, in the limit of vanishing localized potential $V(\Delta)=0$, the system can be mapped onto a nonlinear $\sigma$ model, in which no entanglement phase transition occurs [60]. In the limit $\gamma=0$, in the thermodynamic limit, any finite strength of the SP or disorder ($W$) will lead to localization [73, 103, 69, 102], suppressing long-range information propagation, resulting in an area-law state, and no phase transition occurs in the system. However, due to finite-size effects, the localization length in this case exceeds the system size, which causes the $\gamma=0$ region in Fig. 2(a) and Fig. 8 to still exhibit signatures of a phase transition. For the QPP, Anderson localization occurs only when $V>2J$ [90], which is consistent with the results at $\gamma=0$ in Fig. 2(b).
@@ III.2 Phases and phase transitions
We next follow the finite-size flow of the effective central charge $\bar{c}_{\mathrm{eff}}$ under OBCs. For each size $L$ we fit the entanglement profile to Eq. (13) and obtain a single $\bar{c}_{\mathrm{eff}}$ subjected to different measurement intensities and local potentials, as illustrated in Fig. 4. Near criticality, the half-chain entanglement entropy exhibits logarithmic scaling consistent with non-unitary conformal field theory [104, 105]:
@@
where $\ell$ is the subsystem size, $\bar{c}_{\mathrm{eff}}$ is the effective central charge, and $s_{0}$ is a non-universal constant. In the thermodynamic limit, $\bar{c}_{\mathrm{eff}}\to 0$ in the area-law phase, while it remains finite in the logarithmic phase, thus serving as a diagnostic for the transition [93, 106]. As shown in Fig. 4(a,b), when the local potential is fixed at a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ causes $\bar{c}_{\mathrm{eff}}$ curves for different system sizes to gradually intersect, indicating a transition near $\gamma\approx 0.3$. The insets strongly corroborate this behavior: as $\gamma$ increases, the entanglement entropy transitions from rapid growth to saturation, signaling a crossover from log-law to area-law scaling. Similarly, in Fig. 4(c,d), when the measurement strength is fixed at a low value ($\gamma=0.1$), increasing the local potential strength leads to similar crossing behavior among different system sizes, with transition points occurring around $\Delta\approx 0.3$ and $V\approx 1.8$, respectively. The insets in Fig. 4 clearly demonstrate this effect: the entanglement entropy drops from steep growth to a nearly constant value as localization increases, again signifying a transition from log-law to area-law behavior. These results collectively indicate that both measurement strength and local potential can induce a phase transition in the system. In both cases, entanglement growth is suppressed, although the mechanisms and suppression profiles differ between SP and QPP. These observations reveal how both measurement and localization reduce boundary coupling and suppress entanglement growth. The distinct functional behaviors of SP and QPP reinforce the validity of $\bar{c}_{\mathrm{eff}}$ as a transition indicator.
@@ Appendix A Discussion on the Phase Boundary
this is the entanglement entropy growth rate corrected by the SP. Similarly, for a one-dimensional QPP Eq. (1), the single-particle eigenfunctions localize exponentially with localization length $\xi(V)$. Hence the entanglement entropy growth rate is suppressed to